# Searching for the Shortest-wavelength Aromatic Infrared Bands: No Evidence for the Predicted 1.05 m Polycyclic Aromatic Hydrocarbon Feature

### v1

@@ I Introduction
A feature that has been included in models of interstellar dust, but has yet to be observationally detected, is the PAH feature centered at 1.05 m with a full width at half maximum (FWHM) of 0.058 m. This feature is predicted based on laboratory measurements of electronic transitions in PAH cations (Mattioda et al., 2005b). Since ionized PAHs are believed to be common in the interstellar medium (see Weingartner and Draine, 2001, Figure 9), Draine and Li (2007) included the 1.05 m feature in their model of PAH emission. Although the feature has not been detected, subsequent dust models have continued to include it in the ionized PAH absorption cross sections (e.g., Draine et al., 2021; Hensley and Draine, 2023).
@@ V Discussion
Unlike longer-wavelength PAH features observed at infrared wavelengths (e.g., 3.3, 6.2), the 1.05 feature is not the result of the vibrational modes of PAHs. Instead, the 1.05 m feature is believed to originate from the electronic transitions of ionized PAHs, specifically PAH cations (Mattioda et al., 2005a; Draine and Li, 2007). One possible reason for the absence of this feature perhaps lies with the population of ionized PAHs. The absence of the 1.05 m feature could indicate that the majority of PAHs on this line of sight are neutral. The electronic transitions of neutral PAHs occur at shorter wavelengths (Mattioda et al., 2005b). Indeed, some have hypothesized that these may account for the diffuse interstellar bands (DIBs) observed at optical wavelengths, though such connection remains unconfirmed (Gredel et al., 2011; Salama et al., 2011). While such a scenario would be at odds with the consensus that ionized PAHs are widespread (e.g., Mattila et al., 1996; Onaka et al., 1996), this could suggest more variation in the fraction of ionized versus neutral PAHs than previously assumed.
@@
Regardless, it is clear that our observations are incompatible with current model predictions for the strength of a 1.05 m PAH feature in the interstellar medium.
@@ VI Conclusion
Since the 1.05 m feature is from cationic PAHs, it non-detection poses challenges for our current models of interstellar dust. It suggests revisions may be required to our assumptions on the charge distribution of PAH molecules (i.e., ionized versus neutral, cationic versus anionic) or the types of PAH molecules themselves (i.e., species of PAHs).
@@
Our non-detection of the 1.05 m feature does not preclude the existence of other spectral features from PAH electronic transitions. Since the 1.05 m feature is believed to originate from PAH cations, electronic transitions from neutral PAHs or PAH anions may exist at other near-infrared wavelengths. Additionally, the cationic species that have been measured in the laboratory may not be representative of the PAH cations present in the interstellar medium, which could have electronic transitions at other wavelengths. We nevertheless find no compelling evidence for other interstellar features on the sight line toward BD+40 4223.

### v2

@@ I Introduction
A feature that has been included in models of interstellar dust, but has yet to be observationally detected, is the PAH feature centered at 1.05 m with a full width at half maximum (FWHM) of 0.058 m. This feature is predicted based on laboratory measurements of electronic transitions in PAH cations (Mattioda et al., 2005b). Since ionized PAHs are believed to be common in the interstellar medium (see Weingartner and Draine, 2001b, Figure 9), Draine and Li (2007) included the 1.05 m feature in their model of PAH emission. Although the feature has not been detected, subsequent dust models have continued to include it in the ionized PAH absorption cross sections (e.g., Draine et al., 2021; Hensley and Draine, 2023).
@@ V Discussion
Unlike longer-wavelength PAH features observed at infrared wavelengths (e.g., 3.3, 6.2), the 1.05 m feature is not the result of the vibrational modes of PAHs. Instead, the 1.05 m feature is believed to originate from the electronic transitions of ionized PAHs, specifically PAH cations (Mattioda et al., 2005a; Draine and Li, 2007). One possible reason for the absence of this feature perhaps lies with the population of ionized PAHs. The absence of the 1.05 m feature could indicate that the majority of PAHs on this line of sight are neutral as the electronic transitions of neutral PAHs occur at shorter wavelengths (Mattioda et al., 2005b). However, such a scenario would be at odds with the consensus that ionized PAHs are widespread (e.g., Mattila et al., 1996; Onaka et al., 1996). Dust models have required a certain fraction of PAHs to be ionized in order to replicate observations (e.g., Hudgins and Allamandola, 2004). Should there be more variation in the fraction of ionized versus neutral PAHs than previously assumed, revisions to existing dust models may be required.
@@
Though no prior work has been dedicated to detecting electronic transitions from ionized PAHs (such as the 1.05 m feature), there have been studies to measure such transitions from neutral PAH molecules. In all cases, these efforts have yielded strong upper limits on the presence of electronic transitions. Clayton et al. (2003) studied the UV extinction curve using HST/STIS and found no evidence of absorption features from PAHs. Searches for the electronic transitions of specific neutral PAH molecules (e.g., pyrene, acenaphthene, anthracene) using ground based near-UV and optical spectra also resulted in non-detections (Salama et al., 2011; Gredel et al., 2011). Given the lack of electronic transitions from neutral and positively charged PAHs, one implication may be that the PAHs in the interstellar medium predominantly carry negative charge. However, this is not predicted from physical models of PAHs (e.g., Weingartner and Draine, 2001a) that predict small PAHs to be either neutral or cationic and larger PAHs to be predominantly cationic. On the other hand, it may be that none of the PAH molecules evaluated in laboratory experiments that possess these electronic transitions exist in large quantities in the interstellar medium. In both cases, this further suggests the need for additional laboratory experiments to explore the observational signatures of a wider range of charged and neutral hydrocarbons (for a detailed discussion, see, e.g., Kwok, 2022).
@@ VI Conclusion
Since the 1.05 m feature is from cationic PAHs, its non-detection poses challenges for our current models of interstellar dust. Taking our results here together with non-detections of neutral electronic transitions as well as other deviations from the typical PAH models (e.g., Hudgins et al., 2005; Sadjadi et al., 2015; Tokunaga and Bernstein, 2021; Tokunaga et al., 2025), this suggests revisions may be required to our assumptions on the charge distribution of PAH molecules (i.e., ionized versus neutral, cationic versus anionic) or the chemical nature of the molecules themselves.
@@
Our non-detection of the 1.05 m feature does not preclude the existence of other spectral features from PAH electronic transitions. Since the 1.05 m feature is believed to originate from PAH cations, electronic transitions from neutral PAHs or PAH anions may exist at other near-infrared wavelengths (though, despite dedicated searches, neutral electronic transitions have yet to be confirmed). Additionally, the cationic species that have been measured in the laboratory may not be representative of the PAH cations present in the interstellar medium, which could have electronic transitions at other wavelengths. We nevertheless find no compelling evidence for other interstellar features on the sight line toward BD+40 4223.